OpenStudy (anonymous):

Determine the number of grams of NaCl necessary to make up 100 mL of a 0.10 M solution.

OpenStudy (anonymous):

Molarity = No of moles of Solute / Volume of Solution

OpenStudy (anoop27):

n=M*V n=0.10*1/10 n=0.100 n=mass/Molar mass 0.100*(23+35.5)=Ans